A Bibliography of Publications in American

Mathematical Monthly: 1930–1939

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

02 November 2023 Version 1.05

Title word cross-reference

$1.00 [11]. π [7]. absolute [8]. algebraic [9]. algorithm [6]. analog [12]. arccotangent [7]. Bertrand [11]. Bohr [11]. Book [11]. Carnap [11]. Charles [11]. circle [12]. constant [8]. converse [5]. cyclotomic [1]. Dewey [11]. differential [4]. electric [2]. Encyclopedia [11]. Euclid [6]. Fermat [5]. finding [3]. geometry [4]. infinite [3]. John [11]. Kasner [12]. Khintchine [8]. large [6]. method [3]. Modified [10]. Morris [11]. Neurath [11]. Niels [11]. nine [12]. nine-point [12]. Note [8, 9]. number [2]. numbers [9, 6]. numerical [3]. Otto [11]. photo [2]. photo-electric [2]. plane [12]. point [12]. polynomials [1]. pp [11]. Projective [4].

Quasi [1]. Quasi-cyclotomic [1].

relations [7]. Review [11]. Rudolf [11]. Russell [11].

Science [11]. series [10, 3]. sieve [2]. sum [3].

theorem [5]. trigonometric [9].

Unified [11].

vii [11].

W [11].

References

Lehmer:1932:QCP

[1] D. H. Lehmer. Quasi-cyclotomic polynomials. American Mathematical Monthly, 39(??):383–389, ???? 1932. CODEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1933:PEN

[2] D. H. Lehmer. A photo-electric number sieve. American Mathematical Monthly, 40(??):401–406, ???? 1933. CODEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic).

Camp:1933:NMF

[3] C. C. Camp. A new method for finding the numerical sum of an infinite series. American Mathematical Monthly, 40(??): 537–542, ???? 1933. CODEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lane:1933:PDG

[4] E. P. Lane. Projective differential geometry. American Mathematical Monthly, 40(??):568–579, ???? 1933. CODEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1936:CFT

[5] D. H. Lehmer. On the converse of Fermat’s theorem. American Mathematical Monthly, 43(??):347–354, ???? 1936. CO-DEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1938:EAL

[6] D. H. Lehmer. Euclid’s algorithm for large numbers. American Mathematical Monthly, 45(??):227–233, ???? 1938. CO-DEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1938:AR

[7] D. H. Lehmer. On arccotangent relations for π. American Mathematical Monthly, 45(??):657–664, ???? 1938. CODEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1939:NAC

[8] D. H. Lehmer. Note on an absolute constant of Khintchine. American Mathematical Monthly, 46(??):148–152, ???? 1939. CODEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lehmer:1933:NTA

[9] D. H. Lehmer. A note on trigonometric algebraic numbers. American Mathematical Monthly, 40(??):165–166, ???? 1933. CO-DEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bradshaw:1939:MS

[10] J. W. Bradshaw. Modified series. American Mathematical Monthly, 46(??):486– 492, ???? 1939. CODEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic).

Ballantine:1939:BRB

[11] Constance Ballantine. Book review: Encyclopedia and Unified Science, Otto Neurath, Niels Bohr, John Dewey, Bertrand Russell, Rudolf Carnap, Charles W. Morris, 1938, vii + 75 pp $1.00. American Mathematical Monthly, 46(3):162– 163, March 1939. CODEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/
stable/2302467.

DeCicco:1939:ANP

[12] John De Cicco. An analog of the ninepoint circle in the Kasner plane. American Mathematical Monthly, 46(??):627– 634, ???? 1939. CODEN CACMA2. ISSN 0002-9890 (print), 1930-0972 (electronic).